Assume a consumer with the utility function
U = U(X, Y) = (X + 2)(Y + 1) M = 95, Px= 10, and Py= 5
Set up the consumers’ budget constraint
1
Expert's answer
2018-11-09T15:40:09-0500

The stock turnover is calculated as

Stock turnover=(Cost of sales)/(Average stock held)=(570,000)/(124,250)=4.59 times